OpenStudy (anonymous):

The y-intercept for 3x + 2y = 12 is (0,___).

OpenStudy (nurali):

3x + 2y = 12 x=0 3(0)+2y=12 0+2y=12 2y=12 y=6 (x,y)=(0,6)